OpenStudy (anonymous):

Help FAN & MEDAL !!! Evaluate g(x) = -x + 8 for g(-2)

OpenStudy (anonymous):

SO ITS -6 ?

OpenStudy (anonymous):

is this correct ?

OpenStudy (anonymous):

g(-2) =-(-2)+8 g(-2)=10

OpenStudy (anonymous):

so its -5 ? ..

OpenStudy (anonymous):

no it is : 10

OpenStudy (anonymous):

g(-2) means that x=-2 so you will sub -2 instead of x in the equation

OpenStudy (anonymous):

ok I will try 10 .

OpenStudy (nurali):

g(x) = -x + 8 g(-2)=-(-2)+8 g(-2)=2+8 g(-2)=10